OpenStudy (anonymous):

find the solution of 2r-(5-r)=13+2r

OpenStudy (anonymous):

Distribute the negative then add like terms. (:Those are your first steps. Then you isolate r.

OpenStudy (anonymous):

\[ 2r-(5-r)=13+2r\\ r-5=13\\ r=18 \]

OpenStudy (anonymous):

Step 1. Remove common terms by subtracting 2r from both sides; 2r-(5-r)=13+2r -> -(5-r)=13 Step 2. Distribute the inverse; -(5-r)=13 -> r-5=13 Step 3. Isolate r by adding 5 to both sides. r-5=13 -> r=18